# The low multipoles in the Pantheon+SH0ES data

### v1

@@ 5 Conclusions
In this paper we analysed the Pantheon+ data including a dipole, a quadrupole and a monopole perturbation in the luminosity distance which are motivated by the peculiar motion of the supernovae which leads to an angular dependence of the redshift perturbation. We have found that both, the quadrupole and at very low redshift also the monopole are significant and roughly of the same amplitude as the dipole of the bulk velocity. Removing low redshift supernovae from our sample, we even find that while the monopole and the dipole from the bulk motion are no longer detected with high significance, the quadrupole remains non-zero at more than 95% confidence. It is also interesting to note that the eigenvalues of the quadrupole are even significantly increasing with redshift while their errors stay roughly constant. Hence they also become more significant at higher redshift. This trend is understandable as at higher redshifts, where neither the monopole perturbation nor the dipole (bulk velocity) are significant, the peculiar velocity of the sources has to be modelled by the quadrupole alone in our approach.
Refer to captionRefer to caption
@@
It is intriguing that our result for the bulk velocity is within 95% confidence in agreement with the bulk flow found in the CosmicFlows4 analysis [24]. That it does not perfectly agree with CosmicFlows4 is not surprising since our $z_{\rm cut}$ means that we exclude all supernovae inside a ball of radius $r(z_{\rm cut})$. While the remaining bulk flow is dominated by the one in a shell close to $r(z_{\rm cut})$, this is not quite the same as the velocity inside the ball. Also, while the CosmicFlows4 analysis does include the Pantheon+ data, it has a much larger catalog of about 38’000 galaxy velocities. For $z_{\rm cut}<0.0375$ we obtain a bulk flow of about $316\pm 200$km/s, and based on Appendix B, see also Fig. 12, we find the probability to find a velocity of this size or larger inside a ball of redshift $z_{\rm cut}=0.0375$ is
$P(v^{\rm(bulk)}\geq 316,z_{\rm cut}=0.0375)=8\%\,.$ (5.1)
This is the result within standard $\Lambda$CDM with $\sigma_{kV}(r=112h^{-1}$pc)212$\simeq$ 212) ≃ 212km/s. This seems to be in tension with the standard model at nearly two $\sigma$. However, taking into account that our error in $v^{\rm(bulk)}$ is relatively large we also have to consider $v^{\rm(bulk)}=(316-200)$km/s, for which we find
$P(v^{\rm(bulk)}\geq 116,z_{\rm cut}<0.0375)=82.45\%\,,$ (5.2)
for which certainly there is no tension. The reason that our results are not in strong tension with $\Lambda$CDM is mainly due to the weaker statistical power, hence to the large error bars of $v^{\rm(bulk)}$ and to the fact that we have no truly significant bulk flow at $r\geq 200h^{-1}$pc. Would we take at face value the bulk flow of 280km/s at $z_{\rm cut}=0.1$ which corresponds to $r=300h^{-1}$pc, with $\sigma_{kV}(r=300h^{-1}$pc)98.69$\simeq$ 98.69) ≃ 98.69km/s, we would find
$P(v^{\rm(bulk)}\geq 280,z_{\rm cut}<0.1)=2.3\times 10^{-5}\,.$ (5.3)
This would be similar to the findings of Ref. [24]. But since this bulk velocity is compatible at 68% confidence, see Fig. 14, with the nominally expected value of $\sqrt{3}\times 100$km/s, this is certainly not a valid conclusion.
Within the statistical power of the Pantheon+ data, we therefore conclude that the inferred mmonopole perturbation, dipole and quadrupole in the Pantheon+ data are in reasonable agreement with a velocity field expected in the standard $\Lambda$CDM model of cosmology. It will certainly be very important to repeat this analysis with a larger sample of supernovae. Especially low redshift supernovae with $z\leq 0.1$ are suited to improve the statistical power. For example, the Vera Rubin Observatory’s LSST should be able to detect up to $10^{4}$ supernovae within a year or so in this redshift range [40], while the Pantheon+ dataset has 800 sources in this redshift range.
@@ Acknowledgments
The authors acknowledge financial support from the Swiss National Science Foundation. The computations were performed at University of Geneva using Baobab HPC service.
@@ Appendix A Pantheon+ redshift dependence
In this appendix we also show the contour plots for higher redshift cuts for both the bulk + quadrupole analysis and the full bulk + quadrupole + monopole analysis. We choose $z_{\rm cut}=0.0175,~{}0.025$ and $0.0375$ in Fig. (13, 15) for both the analyses, whereas we set $z_{\rm cut}=0.05$ and $0.1$ in Figs. 14, for only the bulk + quadrupole analysis.
@@ Appendix B Statistical properties of the velocity field
where $f_{0}=f(z=0)$ is the growth factor today in the $\Lambda$CDM model. We show $\sigma_{kV}(R)$ as a function of $R$ in Fig. 16.222For computing the power spectrum used in the probability analysis we used the boltzmann solver camb [41], assuming $H_{0}$=73.4 km/s/Mpc and $\Omega_{{\rm m}}$=0.338 for the background cosmology.
@@
where $\Gamma(k,x)$ is the incomplete Gamma function. We show this probability for $n=3$ in Fig. 17. From Fig. 16 we see that on a scale of $R=112h^{-1}$pc which corresponds to $z_{\rm cut}=0.0375$, we find $\sigma_{kV}\simeq 212$ km/s. In our analysis we find a mean value of $v^{\rm(bulk)}\approx 380$ km/s inside this redshift cut. For a $\chi^{2}(3)$ distribution, we obtain
$P(Z^{2}>3(316/212)^{2})\approx 0.0818\,.$ (B.9)

### v2

@@ 6 Conclusions
In this paper we analysed the Pantheon+ data including a dipole, a quadrupole and a monopole perturbation in the luminosity distance which are motivated by the peculiar motion of the supernovae which leads to an angular dependence of the redshift perturbation. We have found that both the quadrupole and at very low redshift also the monopole are significant and of a comparable amplitude as the dipole of the bulk velocity. Removing low redshift supernovae from our sample, we even find that the monopole and the dipole from the bulk motion are no longer detected with high significance. This can be due to the fact that at higher redshifts, monopole and dipole perturbations, which are sensitive to fluctuations that are at least of the size of the redshift shell, would require fluctuations on ever larger scales which, within standard cosmology are small. However, the quadrupole, that measures fluctuations of typically half the size of the shell remains non-zero at more than 95% confidence. It is also interesting to note that the eigenvalues of the quadrupole are significantly increasing with redshift while their errors stay roughly constant. Hence they also become more significant at higher redshift. This trend is understandable as at higher redshifts, where neither the monopole perturbation nor the dipole (bulk velocity) are significant, the peculiar velocity of the sources has to be modelled by the quadrupole alone in our approach.
@@
It is intriguing that our result for the bulk velocity is within 95% confidence in agreement with the bulk flow found in the CosmicFlows4 analysis [25]. That it does not perfectly agree with CosmicFlows4 is not surprising since our $z_{\rm cut}$ means that we exclude all supernovae inside a ball of radius $r(z_{\rm cut})$. While the remaining bulk flow is dominated by the one in a shell close to $r(z_{\rm cut})$, this is not quite the same as the velocity inside the ball. Also, while the CosmicFlows4 analysis does include the Pantheon+ data, it has a much larger catalog of about 38,000 galaxy velocities. For $z_{\rm cut}<0.0375$ we obtain a bulk flow of about $316\pm 200$km/s, and based on Appendix B, see also figure 16, we conclude that the probability to find a velocity of this size or larger inside a ball of redshift $z_{\rm cut}=0.0375$ is
$P(v^{\rm(bulk)}\geq 316,z_{\rm cut}=0.0375)=16\%\,,$ (6.1)
where we have used the cosmological parameters inferred from the MCMC analysis reported in table 5. This is the result within standard $\Lambda$CDM with $\sigma_{kV}(r=112h^{-1}$pc)212$\simeq$ 212) ≃ 212km/s, and arguably does not appear to be highly unlikely. Taking additionally into account that our error in $v^{\rm(bulk)}$ is relatively large, our analysis does not exclude $v^{\rm(bulk)}=(316-200)$km/s, for which we find
$P(v^{\rm(bulk)}\geq 116,z_{\rm cut}<0.0375)=93.73\%\,,$ (6.2)
and for which certainly there is no tension.
The reason that our results are not in strong tension with $\Lambda$CDM is mainly due to the weaker statistical power of the supernova-only sample, hence to the large error bars of $v^{\rm(bulk)}$ and to the fact that we have no truly significant bulk flow at $r\geq 200h^{-1}$pc. Were we to take at face value the bulk flow of 274km/s at $z_{\rm cut}=0.1$ reported in table 5, which corresponds to $r=300h^{-1}$pc, with $\sigma_{kV}(r=300h^{-1}$pc)98.69$\simeq$ 98.69) ≃ 98.69km/s, we would find
$P(v^{\rm(bulk)}\geq 274,z_{\rm cut}<0.1)=3.82\times 10^{-5}\,.$ (6.3)
[…]
Within the statistical power of the Pantheon+ data, we therefore conclude that the inferred monopole perturbation, dipole and quadrupole in the Pantheon+ data are in reasonable agreement with a velocity field expected in the standard $\Lambda$CDM model of cosmology. It will certainly be very important to repeat this analysis with a larger sample of supernovae. Especially low redshift supernovae with $z\leq 0.1$ are suited to improve the statistical power. For example, the Vera Rubin Observatory’s LSST should be able to detect up to $10^{4}$ supernovae within a year or so in this redshift range [43], while the Pantheon+ dataset has 800 sources in this redshift range.
@@ Acknowledgments
We thank Eoin Ó Colgáin and Subir Sarkar for interesting comments. The authors acknowledge financial support from the Swiss National Science Foundation. The computations were performed at University of Geneva using Baobab HPC service.
@@ Appendix A Pantheon+ redshift dependence
In this appendix we also show the contour plots for higher redshift cuts for both the bulk + quadrupole analysis and the full bulk + quadrupole + monopole analysis. We choose $z_{\rm cut}=0.0175,~{}0.025$ and $0.0375$ in figure (17, 19) for both the analyses, whereas we set $z_{\rm cut}=0.05$ and $0.1$ in figure 18, for only the bulk + quadrupole analysis.
@@
Figure 20: Parameters for redshift cuts at relatively high redshifts for the full bulk + quadrupole + monopole analysis.
@@ Appendix B Statistical properties of the velocity field
where $f_{0}=f(z=0)$ is the growth factor today in the $\Lambda$CDM model. We show $\sigma_{kV}(R)$ as a function of $R$ in figure 21.555For computing the power spectrum used in the probability analysis we used the boltzmann solver camb [44], assuming $H_{0}$=73.6 km/s/Mpc and $\Omega_{{\rm m}}$=0.334 for the background cosmology.
@@
where $\Gamma(k,x)$ is the incomplete Gamma function. We show this probability for $n=3$ in figure 22. From figure 21 we see that on a scale of $R=112h^{-1}$pc which corresponds to $z_{\rm cut}=0.0375$, we find $\sigma_{kV}\simeq 212$ km/s. In our analysis we find a mean value of $v^{\rm(bulk)}\approx 374$ km/s inside this redshift cut. For a $\chi^{2}(3)$ distribution, we obtain
$P(Z^{2}>3(374/212)^{2})\approx 0.01599\,.$ (B.9)